# GCondenser: Benchmarking Graph Condensation

### v2

@@ 4.3 Overall Results (RQ1)
Experimental results show that all GC methods can achieve the highest or second-highest test accuracy in at least one setting. The trajectory-matching method, SFGC, shows a better performance on larger datasets with relatively small budgets. Specifically, SFGC has the best performance on the Arxiv dataset, which is the largest transductive dataset in the experiments. Surprisingly, even the simplest DM methods can effectively condense the graph. In addition, our findings are summarised as follows:
@@
Figure 2: Test accuracy against condensation time for different GC methods on a 90-node condensed graph from the Arxiv dataset, with backbone models GCN and SGC.
@@ 4.7 Cross-architecture Transferability (RQ5)
Figure 3: Transferability of condensed graphs for Cora with budget 35.
@@ 4.8 Continual Graph Learning (RQ6)
To explore the effectiveness of different graph condensation methods in downstream tasks, continual graph learning (CGL) is employed for evaluating the condensed graphs. Three representative methods are selected for comparison: GCond for multi-step matching with structure learning, DosCond for one-step matching with structures learning, and DM for one-step matching without structure learning. These three methods are compared against the random subgraph method, the k-Center graph method, and the whole graph upper-bound. It is worth noting that SFGC, as an offline algorithm, is not suitable to apply in continual graph learning, where online updates are required. The Condense and Train (CaT) framework [19] under the class incremental learning (class-IL) setting is applied. Figure 4 reports the CGL average performance (AP) for CiteSeer and Arxiv datasets. First, the incoming graph is condensed into a 10-node synthetic graph and stored in a memory bank. Then, a GNN model is continually trained only using this memory bank. Hyperparameter choices of baseline methods are presented in Table 8 of Appendix A.5. The results show that using the memory bank generated by DM for CGL can match the performance of using the whole graph. On the contrary, GCond and GCondX lead to a noticeable decline in performance, even incomparable to the random subgraph and the k-Center methods in the Arxiv dataset.
Figure 4: Average performance for CiteSeer (left, three tasks) and Arxiv (right, twenty tasks) under continual graph learning for different graph condensation methods.
@@ 6 Conclusion and Future Work
Limitation and future work: The current state of GCondenser reflects the development of graph condensation methods on datasets with a relatively small scale. In the near future, GCondenser will be expanded to support more emerging GC methods, such as the recently accepted GC-SNTK [26], GEOM [32] and GDEM [18], and additional large-scale graph datasets, such Papers100M [12]. Further, GCondenser will include a wider collection of downstream applications of graph condensation, such as graph neural architecture search as in GraphGym [31], link prediction and graph classification. The straightforward extension to new methods, tasks and datasets will ensure the GCondenser to be an updated and powerful benchmark for graph condensation.
@@ Appendix D Ethic Statement
Appendix E Supplementary
E.1 Support for Newly Proposed GC Methods
GCondenser supports newly proposed methods as well, such as GDEM [18] and GEOM [32]. Due to time constraints, only selected datasets and a limited hyperparameter search are conducted.
Table 10 shows the GDEM condensation performance on relatively smaller graph datasets due to the quadratic time complexity of the Laplacian adjacency matrix decomposition. Table 10 shows the GDEM condensation performance on all baseline datasets, with only the GCN backbone utilised due to the time limit.
The results indicate that GDEM and GEOM have the potential to outperform other baseline methods on specific datasets and condensation sizes. However, due to the novelty of these methods and the large number of hyperparameters involved, a thorough parameter search is not feasible to conduct in the limited time available.
Table 9: Results for GDEM method.
Dataset Size Ratio GDEM
CiteSeer 30 0.9% 71.7 ± 0.3
60 1.8% 72.7 ± 0.6
120 3.6% 73.4 ± 0.4
Cora 35 1.3% 68.0 ± 0.1
70 2.6% 72.8 ± 0.8
140 5.2% 77.4 ± 0.6
PubMed 15 0.08% 73.3 ± 0.6
30 0.15% 78.3 ± 0.8
60 0.3% 78.8 ± 0.3
Table 10: Results for GEOM method.
Dataset Backbone Size Ratio GEOM
CiteSeer GCN 30 0.9% 69.6 ± 0.6
GCN 60 1.8% 67.5 ± 0.9
GCN 120 3.6% 72.1 ± 1.0
Cora GCN 35 1.3% 80.3 ± 1.1
GCN 70 2.6% 81.5 ± 0.8
GCN 140 5.2% 82.2 ± 0.4
PubMed GCN 15 0.08% 80.1 ± 0.3
GCN 30 0.15% 79.7 ± 0.3
GCN 60 0.3% 79.5 ± 0.4
Arxiv GCN 90 0.08% 61.9 ± 0.5
GCN 454 0.15% 65.8 ± 0.4
GCN 909 0.3% 66.2 ± 0.5
Products GCN 612 0.025% 68.5 ± 0.3
GCN 1225 0.05% 69.8 ± 0.3
GCN 2449 0.1% 71.1 ± 0.3
Flickr GCN 44 0.1% 44.6 ± 0.5
GCN 223 0.5% 45.2 ± 0.9
GCN 446 1% 45.5 ± 0.1
Reddit GCN 153 0.05% 90.0 ± 0.5
GCN 769 0.1% 89.4 ± 0.5
GCN 1539 0.2% 91.2 ± 0.1
E.2 More Experiments of Continual Graph Learning
CiteSeer, Cora, Arxiv, Products, Flickr, and Reddit datasets are divided into a series of incoming subgraphs, each with two new classes. If a subgraph contains only one class, this subgraph should be removed from the streaming data. PubMed is not evaluated as it only contains three classes. Offline methods, such as SFGC, GDEM, and GEOM, are also not evaluated. Due to the running time of SGDD, a large number of tasks would be time-consuming. Therefore, SGDD is not evaluated in the CGL.
Table 20 shows that in smaller datasets like CiteSeer and Cora, distribution-matching methods perform better due to the good adaptability of the condensed graphs to the continually learned model. DosCondX can even match the performance of training on the whole graph in long series tasks. However, most methods failed on the Flickr dataset, indicating that this dataset poses challenges for maintaining historical knowledge on the condensed graphs in the CGL setting.

### v3

@@ 4.3 Overall Results (RQ1)
Experimental results show that all GC methods can achieve the highest or second-highest test accuracy in at least one setting. The trajectory-matching methods, SFGC and GEOM, show better performance on larger datasets with relatively small budgets. Specifically, SFGC has the best performance on the Arxiv dataset, while GDEM demonstrates superiority on the Products dataset. Surprisingly, even the simplest DM methods can effectively condense the graph. In addition, although GDEM performs well on small datasets, it encounters scalability issues with large graph datasets. Our findings are summarised as follows:
@@
Figure 2: Test accuracy against condensation time for different GC methods on a 90-node condensed graph from the Arxiv dataset, with backbone models GCN and SGC. More results in Appendix B.1.
@@ 4.7 Cross-architecture Transferability (RQ5)
Figure 3: Transferability of condensed graphs for Cora with budget 35. More results in Appendix B.2.
@@ 4.8 Continual Graph Learning (RQ6)
To explore the effectiveness of different graph condensation methods in downstream tasks, continual graph learning (CGL) is employed for evaluating the condensed graphs. Three representative methods are selected for comparison: GCond for multi-step matching with structure learning, DosCond for one-step matching with structures learning, and DM for one-step matching without structure learning. These three methods are compared against the random subgraph method, the k-Center graph method, and the whole graph upper-bound. It is worth noting that SFGC, as an offline algorithm, is not suitable to apply in continual graph learning, where online updates are required. The Condense and Train (CaT) framework [19] under the class incremental learning (class-IL) setting is applied. CiteSeer, Cora, Arxiv, Products, Flickr, and Reddit datasets are divided into a series of incoming subgraphs, each with two new classes. If a subgraph contains only one class, this subgraph should be removed from the streaming data. PubMed is not evaluated as it only contains three classes. Offline methods, such as SFGC, GDEM, and GEOM, are also not evaluated. Due to the running time of SGDD, a large number of tasks would be time-consuming. Therefore, SGDD is not evaluated in the CGL.
Figure 4 shows that in smaller datasets like CiteSeer and Cora, distribution-matching methods perform better due to the good adaptability of the condensed graphs to the continually learned model. DosCondX can even match the performance of training on the whole graph in long series tasks. However, most methods failed on the Flickr dataset, indicating that this dataset poses challenges for maintaining historical knowledge on the condensed graphs in the CGL setting.
@@ 6 Conclusion and Future Work
Limitation and future work: The current state of GCondenser reflects the development of graph condensation methods on datasets with a relatively small scale. In the near future, GCondenser will be expanded to support more emerging GC methods and additional large-scale graph datasets, such Papers100M [12]. Further, GCondenser will include a wider collection of downstream applications of graph condensation, such as graph neural architecture search as in GraphGym [31], link prediction and graph classification. The straightforward extension to new methods, tasks and datasets will ensure the GCondenser to be an updated and powerful benchmark for graph condensation.